Problem: A supplement of an angle is seven times a complement of the angle. Find the measures of the angle, its complement, and its supplement.
If possible, will you also please show me how you set up the equation to solve this because I would like to learn how to do problems similar to this one as well.

Answer by cleomenius(959)

180 - x = 7(90 - x)
180 - x = 630 - 7x
6x = 450
x = 75 Degrees, the angle.
15 degrees the complement.
105 degrees, the supplement.
7(15) = 105 this does check.
Cleomenius.